# Experience-based Knowledge Correction for Robust Planning in Minecraft

### v2

@@ Abstract
Large Language Model (LLM)-based planning has advanced embodied agents in long-horizon environments such as Minecraft, where acquiring latent knowledge of goal (or item) dependencies and feasible actions is critical. However, LLMs often begin with flawed priors and fail to correct them through prompting, even with feedback. We present XENON (eXpErience-based kNOwledge correctioN), an agent that algorithmically revises knowledge from experience, enabling robustness to flawed priors and sparse binary feedback. XENON integrates two mechanisms: Adaptive Dependency Graph, which corrects item dependencies using past successes, and Failure-aware Action Memory, which corrects action knowledge using past failures. Together, these components allow XENON to acquire complex dependencies despite limited guidance. Experiments across multiple Minecraft benchmarks show that XENON outperforms prior agents in both knowledge learning and long-horizon planning. Remarkably, with only a 7B open-weight LLM, XENON surpasses agents that rely on much larger proprietary models. Code available at https://sjlee-me.github.io/XENON
@@ 4.3 Additional technique: context-aware reprompting (CRe) for controller
In real-world-like environments, an imperfect low-level controller can stall (e.g., in deep water). To address this, XENON employs context-aware reprompting (CRe) technique, which uses an LLM that takes the current image observation and the controller’s language subgoal for item acquisition as input, and outputs whether to replace that subgoal and, if so, a new temporary subgoal to escape the stalled state (e.g., “get out of the water”). Our CRe is adapted from Optimus-1 (Li et al., 2024b) to be suitable for smaller LLMs, with two key differences: (1) a two-stage reasoning process that separates image observation captioning from the subsequent text-only decision to replace the subgoal or not, and (2) a conditional trigger that activates only when the subgoal for item acquisition makes no progress, rather than at fixed intervals. See Appendix˜H for details.
@@ 5.4 Robust dependency learning against knowledge conflicts
Figure 6: Robustness against knowledge conflicts. EGA after 3,000 environment steps in MC-TextWorld under different perturbations of the ground-truth rules. The plots show performance with increasing intensities of perturbation applied to: (a) requirements only, (b) operations only, and (c) both (see Table˜4).
@@
Figure˜6 shows XENON’s strong robustness to knowledge conflicts, as it consistently maintains a near-perfect EGA ($\approx$0.97). In contrast, the performance of all baselines degrades as perturbation intensity increases across all three perturbation scenarios (required items, actions, or both). Notably, we find that prompting an LLM to self-correct is ineffective when the ground truth conflicts with its parametric knowledge: SC shows no significant advantage over DECKARD, which lacks a correction mechanism. ADAM is vulnerable to action perturbations; its strategy of gathering all resource items before attempting a new item fails when the valid actions for those resources are perturbed, effectively halting its learning.
@@ Appendix C Experienced requirement set and dependency graph update
When the agent obtains item $v$ while executing a subgoal $(op,q,u)$, it determines the experienced requirement set $\mathcal{R}_{exp}(v)$ differently depending on whether the operation $op$ is “mine” or falls under “craft” or “smelt”. If $op$ is “mine”, the agent determines $\mathcal{R}_{exp}(v)$ based on the pickaxe in its inventory. If no pickaxe is held, $\mathcal{R}_{exp}(v)$ is $\emptyset$. Otherwise, $\mathcal{R}_{exp}(v)$ becomes $\{(\text{the highest-tier pickaxe the agent has},1)\}$, where the highest-tier pickaxe is determined following the hierarchy: “wooden_pickaxe”, “stone_pickaxe”, “iron_pickaxe”, “diamond_pickaxe”. If $op$ is “craft” or “smelt”, the agent determines the used items and their quantities as $\mathcal{R}_{exp}(v)$ by observing inventory changes when crafting or smelting $v$.
@@ Appendix D Full procedure of XENON
input : invalid operation threshold $x_{0}$, inadmissible item threshold $c_{0}$, less-explored item scale $\alpha_{s}$, inadmissible item scale $\alpha_{i}$
@@ 4   Get an empty inventory invinv
6      Plan $P\leftarrow((a_{l},q_{l},u_{l}))_{l=1}^{L_{v_{g}}}$ by selecting $a_{l}$ for each $u_{l}$, using LLM, $M$, $x_{0}$
@@ 8         Execute $(a,q,u)$ then get the execution result successsuccess
9         Update memory $S(a,u)\mathrel{+}=success$, $F(a,u)\mathrel{+}=\neg success$
@@
12               Reset memory $M[u,\cdot]\leftarrow(0,0)$
@@ 13             $v_{g}\leftarrow\texttt{SelectGoalWithDifficulty}(\hat{\mathcal{G}},C[\cdot])$
Appendix E Details in Adaptive Dependency Learning (ADG)
@@ Appendix F Step-by-step planning using FAM
Given a sequence of aggregated requirements $((q_{l},v_{l}))_{l=1}^{L}$, XENON employs a step-by-step planning approach, iteratively selecting an operation $op_{l}$ for each requirement item $v_{l}$ to make a subgoal $(op_{l},q_{l},v_{l})$. This process considers the past attempts to obtain $v_{l}$ using specific operations. Specifically, for a given item $v_{l}$, if FAM has an empirically valid operation, XENON reuses it without prompting the LLM. Otherwise, XENON prompts the LLM to select an operation, leveraging information from (i) valid operations for items semantically similar to $v_{l}$, (ii) empirically invalid operations for $v_{l}$.
The pseudocode for this operation selection process is detailed in Algorithm˜4. The prompt is shown in Figure˜16.
@@
input : A requirement item $v$, Fine-grained Failure-aware Operation Memory $M[\cdot,\cdot]$, invalid operation threshold $x_{0}$
output : Selected operation $op_{selected}$
@@
3Retrieve history for item $v$: $M_{v}\leftarrow\{(op,n^{succ}_{v,op},n^{fail}_{v,op})\}_{op\in\mathcal{OP}}$
4
/* Find empirically valid and invalid operations for $v$ according to $M_{v}\!\!\!\!\!$ */
5 $\mathcal{OP}^{valid}_{v}\leftarrow\{op\in\mathcal{OP}\mid n^{succ}_{v,op}>0\land n^{succ}_{v,op}-n^{fail}_{v,op}>-x_{0}\}$
6 $\mathcal{OP}^{invalid}_{v}\leftarrow\{op\in\mathcal{OP}\mid n^{succ}_{v,op}-n^{fail}_{v,op}\leq-x_{0}\}$
7
8if $\mathcal{OP}^{valid}_{v}\neq\emptyset$ then
 /* Empirically valid operation found (at most one), reuse it */
9   Select $op_{selected}$ as the single element in $\mathcal{OP}^{valid}_{v}$
10 return $op_{selected}$
@@ 9   Select $op_{selected}$ as the single element in $\mathcal{OP}^{valid}_{v}$
12else
 /* Otherwise, prompt LLM */
13
 /* 1. Collect (item, valid operation) pairs into ItemValidOpPairs */
14   ItemValidOpPairs $\leftarrow\emptyset$
15 foreach $u\in\text{keys}(M)$ do
16      Retrieve history $M_{u}\leftarrow\{(op^{\prime},n^{succ}_{u,op^{\prime}},n^{fail}_{u,op^{\prime}})\}_{op^{\prime}\in\mathcal{OP}}$
17    foreach $op^{\prime}\in\mathcal{OP}$ such that $n^{succ}_{u,op^{\prime}}>0\land n^{succ}_{u,op^{\prime}}-n^{fail}_{u,op^{\prime}}>-x_{0}$ do
18         Add $(u,op^{\prime})$ to ItemValidOpPairs
19
@@ 18         Add $(u,op^{\prime})$ to ItemValidOpPairs
21
 /* 2. Find K pairs from ItemValidOpPairs whose items are Top-K similar to $v$ */
22 $\mathcal{V}_{with\_valid\_op}\leftarrow\{u\mid(u,op^{\prime})\in\text{ItemValidOpPairs}\}$
23   Find similar items $\mathcal{V}_{K}\subseteq\mathcal{V}_{with\_valid\_op}$ based on Top-K semantic similarity to $v$
24 $\{(u_{k},op_{k})\}_{k=1}^{K}\leftarrow\{(u,op^{\prime})\in\text{ItemValidOpPairs}\mid u\in\mathcal{V}_{K}\}$
25
 /* 3. Make candidate operations for LLM, excluding invalid ones for $v$ */
26 $\mathcal{OP}^{cand}_{v}\leftarrow\mathcal{OP}\setminus\mathcal{OP}^{invalid}_{v}$
27
28 if $\mathcal{OP}^{cand}_{v}=\emptyset$ then
29    $\mathcal{OP}^{cand}_{v}\leftarrow\mathcal{OP}$
30
31 $op_{selected}\leftarrow LLM(v,\{(u_{k},op_{k})\}_{k=1}^{K},\mathcal{OP}^{cand}_{v})$
32 return $op_{selected}$
33
34
Algorithm 4 Step-by-step Planning
@@ I.1 Hyperparameters
Failure threshold for invalid action $x_{0}$ 2-2
@@
  • $x_{0}$ (failure threshold for empirically invalid action): Prevents valid operations from being misclassified as invalid due to accidental failures from an imperfect controller or environmental stochasticity. Values that are too small or large hinder dependency learning and planning by hampering the discovery of valid actions.
@@ I.2 Human-written plans
The original ADAM started with an initial inventory containing 32 quantities of experienced resource items (i.e., items used for crafting other items) and 1 quantity of tool items (e.g., pickaxes, crafting table), implicitly treating those items as a predicted requirement set for each item. Its memory recorded which operations were used for each subgoal item without noting success or failure, and its intrinsic goal selection was guided by an expert-defined exploration curriculum.
In our experiments, ADAM starts with an empty initial inventory. The predicted requirements for each goal item in our ADAM implementation assume a fixed quantity of 8 for all resource items. This quantity was chosen to align with $\alpha_{i}$, the hyperparameter for the quantity scale of requirement items for inadmissible items, thereby ensuring a fair comparison with XENON. The memory stores successful operations for each item, but did not record failures. This modification aligns the memory mechanism with SC and DECKARD baselines, enabling a more consistent comparison across baselines in our experimental setup. Intrinsic goal selection is random, as we do not assume such an expert-defined exploration curriculum.
@@ 14 Craft - Execute crafting to obtain new item
MC-Textworld is a text-based environment based on Minecraft game rules (Zheng et al., 2025). We employ Minecraft version 1.16.5. In this environment, basic rules and goals are the same as those in the MineRL environment Section˜J.2. Furthermore, resource item spawn probabilities do not need to be adjusted, unlike in MineRL Section˜J.2.5. This is because an agent succeeds in mining an item immediately without spatial exploration, if it has a required tool and “mine” is a valid operation for that item.

### v3

@@ Abstract
Large Language Model (LLM)-based planning has advanced embodied agents in long-horizon environments such as Minecraft, where acquiring latent knowledge of goal (or item) dependencies and feasible actions is critical. However, LLMs often begin with flawed priors and fail to correct them through prompting, even with feedback. We present XENON (eXpErience-based kNOwledge correctioN), an agent that algorithmically revises knowledge from experience, enabling robustness to flawed priors and sparse binary feedback. XENON integrates two mechanisms: Adaptive Dependency Graph, which corrects item dependencies using past successes, and Failure-aware Action Memory, which corrects action knowledge using past failures. Together, these components allow XENON to acquire complex dependencies despite limited guidance. Experiments across multiple Minecraft benchmarks show that XENON outperforms prior agents in both knowledge learning and long-horizon planning. Remarkably, with only a 7B open-weight LLM, XENON surpasses agents that rely on much larger proprietary models. Project page: https://sjlee-me.github.io/XENON
@@ 4.3 Additional technique: context-aware reprompting (CRe) for controller
In real-world-like environments, an imperfect controller can stall (e.g., in deep water). To address this, XENON employs context-aware reprompting (CRe), where an LLM uses the current image observation and the controller’s language subgoal to decide whether to replace the subgoal and propose a new temporary subgoal to escape the stalled state (e.g., “get out of the water”). Our CRe is adapted from Optimus-1 (Li et al., 2024b) to be suitable for smaller LLMs, with two differences: (1) a two-stage reasoning process that captions the observation first and then makes a text-only decision on whether to replace the subgoal, and (2) a conditional trigger that activates only when the subgoal for item acquisition makes no progress, rather than at fixed intervals. See Appendix˜H for details.
@@ 5.4 Robust dependency learning against knowledge conflicts
Figure 6: Robustness against knowledge conflicts. EGA after 3,000 environment steps in MC-TextWorld under different perturbations of the ground-truth rules. The plots show performance with increasing intensities of perturbation applied to: (a) requirements only, (b) actions only, and (c) both (see Table˜4).
@@
Figure˜6 shows XENON’s robustness to knowledge conflicts, as it maintains a near-perfect EGA ($\approx$0.97). In contrast, the performance of all baselines degrades as perturbation intensity increases across all three perturbation scenarios (required items, actions, or both). We find that prompting an LLM to self-correct is ineffective when the ground truth conflicts with its parametric knowledge: SC shows no significant advantage over DECKARD, which lacks a correction mechanism. ADAM is vulnerable to action perturbations; its strategy of gathering all resource items before attempting a new item fails when the valid actions for those resources are perturbed, effectively halting its learning.
@@ 5.6 Ablation studies on hyperparameters
We provide additional ablations in the Appendix on dependency and action learning—when initializing the dependency graph from an external source mismatched to the environment (Figure˜23), when scaling to more goals/actions (Figure˜24), and when using a smaller 4B planner LLM (Figure˜26)—as well as an ablation of action selection methods for subgoal construction (Figure˜25).
@@ 6 Conclusion
Acknowledgments
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) and IITP-ITRC (Information Technology Research Center) grant funded by the Korea government (MSIT) (No. RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); IITP-2026-RS-2024-00437866; RS-2024-00509258, Global AI Frontier Lab), by a grant from the Korea Institute for Advancement of Technology (KIAT), funded by the Ministry of Trade, Industry and Energy (MOTIE), Republic of Korea (RS-2025-00564342), and by Seoul R&BD Program (SP240008) through the Seoul Business Agency (SBA) funded by The Seoul Metropolitan Government.
@@ Appendix C Experienced requirement set and dependency graph update
When the agent obtains item $v$ while executing a subgoal $(a,q,u)$, it determines the experienced requirement set $\mathcal{R}_{exp}(v)$ differently depending on whether the high-level action $a$ is “mine” or falls under “craft” or “smelt”. If $a$ is “mine”, the agent determines $\mathcal{R}_{exp}(v)$ based on the pickaxe in its inventory. If no pickaxe is held, $\mathcal{R}_{exp}(v)$ is $\emptyset$. Otherwise, $\mathcal{R}_{exp}(v)$ becomes $\{(\text{the highest-tier pickaxe the agent has},1)\}$, where the highest-tier pickaxe is determined following the hierarchy: “wooden_pickaxe”, “stone_pickaxe”, “iron_pickaxe”, “diamond_pickaxe”. If $a$ is “craft” or “smelt”, the agent determines the used items and their quantities as $\mathcal{R}_{exp}(v)$ by observing inventory changes when crafting or smelting $v$.
@@ Appendix D Full procedure of XENON
input : invalid action threshold $x_{0}$, inadmissible item threshold $c_{0}$, less-explored item scale $\alpha_{s}$, inadmissible item scale $\alpha_{i}$
@@ 7       $v_{g}\leftarrow\texttt{SelectGoalWithDifficulty}(\hat{\mathcal{G}},C[\cdot])$
10      Plan $P\leftarrow((a_{l},q_{l},u_{l}))_{l=1}^{L_{v_{g}}}$ by selecting $a_{l}$ for each $u_{l}$, using LLM, $S$, $F$, $x_{0}$
@@ 12         Execute $(a,q,u)$ then get the execution result successsuccess
13
14       if success then $S(a,u)\leftarrow S(a,u)+1$
15       else $F(a,u)\leftarrow F(a,u)+1$
16
@@
19               Reset memory $S(\cdot,u)\leftarrow 0,F(\cdot,u)\leftarrow 0$
@@ 20             $v_{g}\leftarrow\texttt{SelectGoalWithDifficulty}(\hat{\mathcal{G}},C[\cdot])$
Appendix E Details in Adaptive Dependency Graph (ADG)
@@ Appendix F Step-by-step planning using FAM
Given a sequence of aggregated requirements $((q_{l},v_{l}))_{l=1}^{L}$, XENON employs a step-by-step planning approach, iteratively selecting an high-level action $a_{l}$ for each requirement item $v_{l}$ to make a subgoal $(a_{l},q_{l},v_{l})$. This process considers the past attempts to obtain $v_{l}$ using specific actions. Specifically, for a given item $v_{l}$, if FAM has an empirically valid action, XENON reuses it without prompting the LLM. Otherwise, XENON prompts the LLM to select an action, leveraging information from (i) valid actions for items semantically similar to $v_{l}$, (ii) empirically invalid actions for $v_{l}$.
The pseudocode for this action selection process is detailed in Algorithm˜4. The prompt is shown in Figure˜16.
@@
Input : An item $v$, Action set $\mathcal{A}$, Success/Failure counts from FAM $S(\cdot,\cdot)$ and $F(\cdot,\cdot)$, Invalid action threshold $x_{0}$
Output : Selected action $a_{selected}$
@@
/* 1. Classify actions based on FAM history (S and F counts) */
3 $\mathcal{A}^{valid}_{v}\leftarrow\{a\in\mathcal{A}\mid S(a,v)>0\land S(a,v)>F(a,v)-x_{0}\}$
4 $\mathcal{A}^{invalid}_{v}\leftarrow\{a\in\mathcal{A}\mid F(a,v)\geq S(a,v)+x_{0}\}$
5
6if $\mathcal{A}^{valid}_{v}\neq\emptyset$ then
 /* Reuse the empirically valid action if it exists */
7   Select $a_{selected}$ from $\mathcal{A}^{valid}_{v}$
8 return $a_{selected}$
9
10else
 /* Otherwise, query LLM with similar examples and filtered candidates */
@@ 7   Select $a_{selected}$ from $\mathcal{A}^{valid}_{v}$
 /* (i) Retrieve valid actions from other items for examples */
12 $\mathcal{V}_{source}\leftarrow\{u\in\hat{V}\setminus\{v\}\mid\exists a^{\prime},S(a^{\prime},u)>0\land S(a^{\prime},u)>F(a^{\prime},u)-x_{0}\}$
13   Identify $\mathcal{V}_{topK}\subseteq\mathcal{V}_{source}$ as the $K$ items most similar to $v$ (using S-BERT)
14 $\mathcal{D}_{examples}\leftarrow\{(u,a_{valid})\mid u\in\mathcal{V}_{topK},a_{valid}\in\mathcal{A}^{valid}_{u}\}$
15
 /* (ii) Prune invalid actions to form candidates */
16 $\mathcal{A}^{cand}_{v}\leftarrow\mathcal{A}\setminus\mathcal{A}^{invalid}_{v}$
17
18 if $\mathcal{A}^{cand}_{v}=\emptyset$ then
19    $\mathcal{A}^{cand}_{v}\leftarrow\mathcal{A}$
@@
21 $a_{selected}\leftarrow\text{LLM}(v,\mathcal{D}_{examples},\mathcal{A}^{cand}_{v})$
22 return $a_{selected}$
23
Algorithm 4 Step-by-step Planning with FAM
@@ I.1 Hyperparameters
Failure threshold for invalid action $x_{0}$ 22
@@
  • $x_{0}$ (failure threshold for empirically invalid action): Prevents valid actions from being misclassified as invalid due to accidental failures from an imperfect controller or environmental stochasticity. Values that are too small or large hinder dependency learning and planning by hampering the discovery of valid actions.
@@ I.2 Human-written plans
The original ADAM started with an initial inventory containing 32 quantities of experienced resource items (i.e., items used for crafting other items) and 1 quantity of tool items (e.g., pickaxes, crafting table), implicitly treating those items as a predicted requirement set for each item. Its memory recorded which actions were used for each subgoal item without noting success or failure, and its intrinsic goal selection was guided by an expert-defined exploration curriculum.
In our experiments, ADAM starts with an empty initial inventory. The predicted requirements for each goal item in our ADAM implementation assume a fixed quantity of 8 for all resource items. This quantity was chosen to align with $\alpha_{i}$, the hyperparameter for the quantity scale of requirement items for inadmissible items, thereby ensuring a fair comparison with XENON. The memory stores successful actions for each item, but did not record failures. This modification aligns the memory mechanism with SC and DECKARD baselines, enabling a more consistent comparison across baselines in our experimental setup. Intrinsic goal selection is random, as we do not assume such an expert-defined exploration curriculum.
@@ 14 Craft - Execute crafting to obtain new item
MC-Textworld is a text-based environment based on Minecraft game rules (Zheng et al., 2025). We employ Minecraft version 1.16.5. In this environment, basic rules and goals are the same as those in the MineRL environment Section˜J.2. Furthermore, resource item spawn probabilities do not need to be adjusted, unlike in MineRL Section˜J.2.5. This is because an agent succeeds in mining an item immediately without spatial exploration, if it has a required tool and “mine” is a valid action for that item.